OpenStudy (anonymous):

Find the distance between the points (-4, 3) and (5, -1)

OpenStudy (anonymous):

distance formula\[d=\sqrt{(x_2-x_1)^2 +(y_2-y_1)^2}\]

OpenStudy (anonymous):

I did it and ended up with \[\sqrt{65}\] is that right???

jimthompson5910 (jim_thompson5910):

that's not correct, how did you get 65?

OpenStudy (anonymous):

\[\sqrt{(5+4)^{2}+ (-1-3)^{2}} = \sqrt{9^{2 }- 4^{2}} = \sqrt{81-16}\]

jimthompson5910 (jim_thompson5910):

it should be 81+16 not 81-16

OpenStudy (amistre64):

\[\sqrt{(5+4)^{2}+ (-1-3)^{2}} = \sqrt{9^{2 }- 4^{2}} = \sqrt{81-16}\] ^^ bad :)

OpenStudy (anonymous):

now i got the square root of 162...

OpenStudy (anonymous):

its +16 not -16 because you're squaring a negative number and a negative number square is always positive

OpenStudy (anonymous):

So if it's 81+16 it ends up being the square root of 97? And I just leave it as that?

OpenStudy (anonymous):

well if it cant be factored then you just leave it as it is

OpenStudy (anonymous):

is that what you got too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you!!